2023 August/September Restart OUTLINE

UT September (Dates TBD) - Engineering On-Sky Restart Block

Sept 1-9 Facility instruments, PEPSI checkout, DX LUCI AO

Sept 21-22 SX AO Checkout, LBTI checkout.

Instruments Released for Science

  • PEPSI-PFU - as of
  • LBCs - as of
  • LUCIs - as of
  • MODS- as of
  • DX LUCI AO - as of
  • SX LUCI AO - as of
  • LBTI - as of

Breakdown of Dates

  • Daytime tasks in local time

CLOSED DOME

  • Infrastructure checkout
    • DONE SX M3 Install 8/17
    • DONE DX M3 Install 9/6
    • DONE AGW-1 Install 8/24 + 2 days for check out
    • DONE AGW-2 Already on telescope
    • DONE DX Adsec Install 8/21
      • DONE Initial SL flat check with MODS 8/23 (partial) 8/28
      • DONE DX Seeing limited checkouts
      • DX AO checkouts
    • DONE SX Adsec Install 8/31
      • DONE SX Seeing limited checkouts
      • SX AO calibrations and checkouts
  • Instrument Installs and Closed dome
    • LUCI
      • LUCI2
        • DONE Install 8/17
        • Cooling on telescope 8/21-8/25 - cooling paused after leak found. Warming. To be removed for repairs.
        • DONE LUCI2 leak repair
        • DONECooling in lab
        • DONE Install - Installed 10/16
        • DONE Cold checks 1day -10/18
        • DONE Grating and Wavelength Calibration 6 hours
        • DONE Flexure Compensation 4days (1.2 days for 3 commons modes)
        • DONE Flexure verification 0.5 days
        • DONELUCI2 cabex took place with the instrument open.
      • LUCI1 - delayed install for G040 grating refab. Potential conflicts with AO during recalibration
        • Cooling LUCI1 in lab
        • DONE Install LUCI1 10/26
        • DONE Cold checks 2day - with caveats
        • Grating and Wavelength Calibration 6 hours (
        • Flexure Compensation 4days (1.2 days for 3 commons modes)
        • Flexure verification 0.5 days
        • DONELUCI1 cabex took place with the instrument open.
    • MODS (installed 8/15)
      • DONE Install 8/15
      • DONE Wake-up and Functional Checkout 8/22
      • DONE Cooling MODS 8/24
      • DONE AGW 5 & 6 check 8/23
      • DONE IMCS TTF zeropoints - taken manually as issues with MODS1 blue quadcell persists (9/1)
      • DONE Focus check
      • DONE Mask switch adjustment and Mask exchange 8/27 & 8/28 & 8/29
      • DONE Basic Calibrations for reference 4+hrs 8/30
    • TMS
      • DONE TMS installation (1d). 8/24
      • DONE TMS Alignment (2d) 8/25-26
      • DONE Update TMS matrix
      • DONE TMS remote shutter operability test (0.5hr)
      • DONE TMS Channel Strengths - Moving telescope (4hr)
      • DONE Channel Dropping test
    • LBC (8hrs)
      • DONE Install Aug 15
      • DONE Rotator and FW Checkout 8/16
      • DONE CCD Checkout 8/22-
      • DONE Check filters 1hr closed dome 8/22-
      • DONE Swing arm repeatability test 4hrs 8/27 (tentative)
      • DONE Clean LBC Lenses 8/30 (tentative)
    • PEPSI ()
      • DONE PEPSI PFU AGW7/8 Functional checkout (8/25)
      • DONE Off telescope detector install and red CD repair (8/16-8/30)

ON-SKY

  • Telescope General (1 night)
  • LUCI (17hr)
  • MODS (13.5)
  • LBC (9hr)
  • PEPSI (10 hr)
  • LBTI (Science camera and AO) (2x0.5d)
  • LUCI AO (1d)

General Scope of Activities

  • Closed Dome/Daytime Instrument Checks (each instrument must be released before night-time operations)
  • Telescope Wake-Up
  • Telescope Infrastructure
    • AGW1/2
    • M3 SX/DX
  • Checkout of Facility Instruments
  • LBTI Restart (1-2 hr)
    • Close loop w/WFS only during restart (science cameras may not be cold enough)
    • Close loop w/science camera
  • AO (SX) Restart (2 x 0.5 nights)
    • Daytime calibrations required

Closed Dome/Daytime Work Details

LBC (8 hrs)

LBC checkout is outlined in the document: LBC Restart Plan
  • DONE Rotator Check
    • At least 2 people: ISP up at the LBC and ISA/SO running testrotator root@lbccontrol
      cd /lbccontrol/current
      ./testrotatormove the rotator to a few positions as directed by the ISP who will look for any signs of potential snags or pinching of cables.
  • DONE Filter wheel check
    • May start this immediately after rotator test, but continue after ISP has gone down. root@lbccontrol
      cd /lbccontrol/current
      testfiltersMove filters. Could also measure all overheads in switching filters. These range from ~7 - 20 seconds, depending on which filters you are moving from and to.30 pairs for B and 56 for R
  • DONE CCD Check
    • In conjunction with Alessandro.
      • are bias levels ~300 ADU for all chips?
      • is there any noise (e.g. horizontal banding LBCR)
    • Through Eng level commands, we can take test images without having to run up all of the camera systems (both sci & tec PCs, filter and rotator)
         root@lbccontrol (ssh -X or ssh -Y this time to redirect display)</br>
           cd /lbccontrol/current</br>
    1. use ./power.py to turn on PC & controller
      Press button for PC to turn on
      Press button for controller to turn on
      Refresh
    2. Use the KVM or a VNCViewer to connect to the PC
    3. testcamera -
      must set the controller number
      initialize
      save files as mef
      display with ds -> file -> open -> mosaic image iraf
  • DONE Check Filters Closed Dome

    Light levels: Gallery & stairwells only
    Script name: /lbt/lbto/Calib_OBs/ClosedDome/AllFilt_Bright_PA0.xml
    These use different exposure times, but you can still check for the annealing pattern characteristic at U & B and that the count levels are reasonable

TMS Testing for LBCs (4hrs)

DONE The procedure is just sketched out below…
  1. Power on TMS lasers
  2. As lbto@obs1,2 or robs, or as telescope@robs, bring up TCSGUI and from it, IIF, PSF and OSS GUIs
  3. Configure the telescope for the LBCB and LBCR (usually ask someone on the mountain)
  4. Authorize the LBCs (use IIFGUI -> Control -> IIF Control -> select LBCB and LBCR in the pull-down menus and then click Authorize.
  5. Click collimate on the PSF GUIs
  6. With the LBCs deployed, take a TMS reference on both SX and DX sides:
    1. open an terminal window as user telescope on robs
    2. cd /lbt/lbto/supportscripts/MetrologySupport/
    3. ./lbc_set_tms_offset.py -s 2
  7. repeat at least twice to get a few references
  8. Use the OSS GUI to retract the swing arms and then deploy them again (screenshot can be added…)
  9. Go back to step 5
  10. Repeat steps 6-8 about 9 or 10 times.

MODS (1d)

MODS restart plan is outlined: MODS Restart Plan

The list below comes from the MODS Post-shutdown Startup wiki page

  1. DONE MODS wake-up through coldStart
    1. pingMODS to check all connections
    2. monitor cooldown:
      • Start isis, ie, agw, env services
      • Turn on the HEBs if glycol flow is OK
      • start IC programs
      • turn MODS alerts back on
    3. pokeMODS
    4. Run coldStart scripts (~15 min)
      • Prefer telescope is at or near zenith and not moving.
  2. DONE AGw checkout (~30 min)
    • Needs some dome lights, but maybe dim. And any orientation is fine.
  3. DONE Functional checkout (~30 min)
    • Any lights, any orientation ok
  4. DONE ImSnaps (~10 min)
    • Any lights, any orientation though prefer not horizon.
  5. DONE IMCS TTF zeropoints (~20 min)
    • Telescope @ 60 deg and LDG/RDG = 207/195.
  6. DONE Check focus settings (~20 min per script)
    • Try to get through the dual+red grating and imaging at least and the rest can be done as time permits.
      • Grating DONE dual, DONE red
      • Imaging dual, DONEred
      • Grating blue DONE
      • Imaging blue DONE
      • Prism dual, DONE red, DONE blue
  7. DONE Basic calibrations for reference (3 hrs + )
    • Telescope at zenith, lights off for all but the lamps. Take these as able, but try to get through the (1) BIases, (2) Comparison lamps (grating), (3) Slitless pixel flats and (4) Imaging flats and then the dual grating slit flats.
    • Estimated time for (1)-(4) is about 3 hours. (Estimates from /lbt/lbto/mods/howlong.py)
    1. DONE Biases: 25 min
      • DONE 8K — 9 min
      • DONE 3K — 6 min
      • DONE 4K — 6 min
      • DONE 1K — 4 min
    2. DONE Slitless pixel flats: 51 min
      • DONE dual - 20 min
      • DONE red - 11 min
      • DONE blue - 20 min
    3. DONE Comparison lamps: 25 min
      • DONE dual - 8 min
      • DONE red - 7 min
      • DONE blue - 8 min
    4. DONE Imaging flats: 71 min*
      • DONE dual - 35 min
      • DONE red - 19 min
      • DONE blue - 17 min
1 DONE Slit flats:
      • Dual (113 min, w/o 0.3")
        • DONE 0.6" - 20 min
        • DONE 0.8" - 19 min
        • DONE 1" - 19 min
        • DONE 1.2" - 19 min
        • 2.4" - 18 min
        • 5" - 18 min
        • 0.3" -
      • DONE Red (67, w/o 0.3")
        • DONE 0.6" - 12 min
        • DONE 0.8" - 12 min
        • DONE 1" - 12 min
        • DONE 1.2" - 12 min
        • DONE 2.4" - 12 min
        • DONE 5" - 7 min
        • 0.3"
      • Blue (81, w/o 0.3")
        • DONE 0.6" - 14 min
        • DONE 0.8" - 14 min
        • DONE 1" - 14 min
        • DONE 1.2" - 13 min
        • DONE 2.4" - 13 min
        • DONE 5" - 13 min
        • 0.3" - 15 min
    1. DONE Comparison lamps (prism = prlamps)
      • 23 min total - good check on whether all the lamps are working.
        • dual - 8 min
        • red - 7 min
        • blue - 7 min
  1. DONE Mask Exchange (2hrs)
    1. Follow procedure:https://wiki.lbto.org/Instrumentation/MODSMaskExchange
    2. Install astrometric masks.
    3. Estimated time 2hrs

LUCIs (5-6 days)

  • Basic Checks needed for release for engineering (and AO) work - 1 day in total
    • non-MOS Functional Checkout on telescope L1 (cold) - 0.5 day - internal - no restrictions on telescope
    • DONE non-MOS Functional Checkout on telescope L2 (cold) - 0.5 day- internal - no restrictions on telescope
    • MOS Functional Checkout on telescope L1 (cold) - 0.5 day - internal - no restrictions on telescope
    • DONE MOS Functional Checkout on telescope L2 (cold) - 0.5 day - internal - no restrictions on telescope
    • PFS alignment and Camera and Grating refocus L1 - 1hr - requires L1 authorized and M3 aligned
    • DONE PFS alignment and Camera and Grating refocus L2- 1hr - requires L2 authorized and M3 aligned
  • Grating Wavelength Calibration - 6hr take data, 1-2 days process data (cal unit in beam - stationary telescope)
    • Scripts and README located in:

      /home/lbto/dthompson/LUCI/Calibration/GratingCalibration

    • DONE G210 L2
    • DONE G200 L2
    • DONE G040 L2
    • G200 L1
    • G210 L1
    • G040 L1
  • Flexure Checks - 3 days? - Requires moving telescope (moving instrument rotator and elevation drive) - Starting (Requires basic checks to be completed first)
    • Scripts and README located in:

      /home/lbto/dthompson/LUCI/Calibration/Flexure_Compensation

    • Starting with basic modes of LUCI2
      • DONE G210/N1.8 - L2
      • DONE G200/N1.8 - L2
      • DONE N3.75 imaging - L2
      • DONE N30 Imaging L2
    • Once LUCI1 Installed, complete critical modes (and execute bino any remaining modes)
      • G210/N1.8 - L1
      • G200/N1.8 - L1
      • N3.75 imaging - L1
      • N30 Imaging L1
    • Lower priority modes
      • G210/N3.75 - L1/L2
      • G200/N3.75 - L1/L2
      • G040/N30 - L1/L2
      • N1.8 imaging - L1/L2
    • Verification (0.5days) - 1-2 Common modes are sufficient.
  • DONE Mask Exchange ( Masks will be inserted while instruments are open in the lab, no cabex required)

PEPSI (? day for each focal station)

  • DONE PEPSI PFU AGW7/8 checks that fully functional (homed and exercised) as well all
    four AzCam are readable.
  • DONE Internal funcitonal checkout
    • PFU
    • Calibration

LBTI (2 days)

  • DONEInternal Checkout (0.5 day)
  • DONE AdSec check out (1.5 day)

AO

SX - (5 days seeing limited, +6-21days AO)
  • DONE Install Aug 30th
  • DONE For Seeing limited
    • Measure Feed Forward Matrix after install (1d) (Not Needed_
    • DONE Measure New Flat (4d)
  • DONE For AO
    • If flat good then verify performance with old reconstructors in all observing modes (6d)
    • If flat not good then remeasure new KL modal basis set and new reconstructors (21days)
DX - (3 days seeing limited + 6-21 days for AO)
  • DONE DX installed on Aug 21, but requires tertiary for checkout which will be installed Aug 31
  • DONE For Seeing limited
    • DONE Seeing limited flat verified with ARGOS source and MODS2 8/29
      The reconstructed wavefront errors were
      Raw Zernikes
      Zern # sext cow
      ---------------------------
      Z[ 2] = 0 0
      Z[ 3] = 0 0
      Z[ 4] = 0 -74
      Z[ 5] = 0 112
      Z[ 6] = 0 940
      Z[ 7] = 0 -129
      Z[ 8] = 0 -201
      Z[ 9] = 0 132
      Z[10] = 0 -106
      Z[11] = 0 -27
      The current DX AdSec Flat is (quite!) good enough for all Gregorian Focal Stations
      for Seeing Limited observing when we start on-sky operation on Friday.
    • Further verification of the flat?
  • DONE For AO
    • If flat good then verify performance with old reconstructors in all observing modes (6d)
    • (NOT Needed) If flat not good then remeasure new KL modal basis set and new reconstructors (21days)


Potential Conflicts/Issues


List of On-Sky Tasks

Telescope Wake-up (1 hr)

  • * Closed Dome Test Presets (multiple instruments)
    • DONE LBC
    • LUCIs
    • DONE PEPSI PFU
    • DONE MODS
  • On-sky presets (Track, Acquire, Active)
    • DONE LBC Track
    • LUCIs Active, Acquire, Track
    • DONE PEPSI PFU - Active, Acquire
    • DONE MODS Active, Acquire, Track

Telescope Infrastructure (Critical)

Overview
  1. AGw1/2 Alignment (Priority critical)
  2. M3 Alignment LFBG, RFBG (Priority critical)
  3. Rotator Center LFBG, RFBG
  4. Field Aberration (Priority critical)
  5. DONE AGw7/8 Alignment (Priority critical)
  6. DONE M3 alignment PFU (Priority critical)
  7. DONE AGw5/6 Alignment (Priority critical)
  8. DONE Rotator Center MODS (Priority critical)
  9. DONE Field Aberration LDG, RDG (Priority dependent #3)
Details
  1. AGW1/2 Alignment (LUCI1/2)
  2. M3 Alignment LFBG,RFBG
  3. Rotator Center on LUCI detectors (Field center)
    • Scripts located in LUCI_2022_Restart_V1.83.xml under Field Center or for M52 generated in Service_Observing/Restart_2023/LUCI/Rot_center
    • LUCI configured: N30 fieldstop aligned, N375 camera, mirror, Ks
      • Run the sequence:
        • active 0deg to collimate, then scripts for
        • track 0deg
        • track 180deg
        • track 90deg
        • track 270deg
  4. Field Aberration
  5. DONE AGw7/8 Alignment
    • Verify SH
  6. DONE M3 Alignment PFU
    • Pupil alignment check - find star and center it
  7. DONE AGw5/6 Alignment (MODS1/2)
    • Verify SH spots
  8. DONE Rotator Centers on MODS detectors
    • Scripts and readme located in: /home/lbto/Service_Observing/Restart_2023/MODS/RotCen
    • ~30 minutes. 10 for the initial slew+collimation and then about 5 min per PA (~3 for the script, but add time for rotator slews)
    • Run the sequence:
      • active0 to collimate, then
      • track0
      • track180
      • track90
      • track270
  9. DONE Field Aberration LDG, RDG

LBC (high)

Overview
  1. Twilight flats - ~30-min per morn/eve twilight periods (Priority high)
  2. DONE Check hub positions ~ 10-min (Priority high)
  3. DONE collimate & copoint ~30 min (Priority high)
  4. DONE rotator center ~15 min (Priority high)
  5. DONE focal plane tilt (~30 min for collimate+PA0,180,0 set) (Priority high)
  6. DONE observe with guiding & dithering ~15-min (Priority med)
  7. DONE observe NS field ~30-min minimum (Priority low)
  8. DONE photometric standards (1 hr, ~30 min/field and 2 fields) (Priority high)
  9. DONE TMS testing, in the background for 5-7. (Priority high)
  10. Pointing model? (Priority low)
Details
  1. Twilight flats: Try to obtain flats through all 6 B and 8 R filters during the course of restart
    • LBCB *U-Bessel
      • V-Bessel
      • SDT_USpec
      • B-Bessel
      • r-Sloan
      • g-Sloan
    • LBCR
      • r-Sloan
      • I-Bessel
      • i--Sloan
      • z-Sloan
      • R-Bessel
      • V-Bessel
      • F972N20
      • Y-Fan
  2. DONE Check LBCR hub positions.
    1. Use a copointing script to slew to a field at ~70-75 deg elevation.
    2. Use /lbt/lbto/lbc/Calib_OBs/FOCUS/ Choose either RB_rVfastextra4.ob which will use all 4 chips or RB_rVlowextra.ob - uses the same portion of chip 2 as dofpia
    3. Measure diameters of extra-focal images for B and R. Compare. Ideal diameter is 61.5 pix, though temperature will affect this, so compare the red with the blue pupils. The blue winter/spring/summer hub settings are fine. What focus offset (call it delZ_R) is needed to make red match blue? Scaling factors are: https://scienceops.lbto.org/lbc/observing-with-lbc/troubleshooting/#TOC-collimation-problems: Increasing z4 or Z will decrease the diameter of the pupil. delta_Z[mm]/delta_d[pix] = 1/75.32 = 0.013 mm/pix
    4. Adjust PSF -> Z global offset by delZ_R
    5. Repeat OB
    6. Let the ISp know that this offset is needed to LBCR summer/spring/winter positions.
  3. DONE Pointing and Collimation
    1. Collimate Use the same copointing script to collimate
    2. Correct pointing/copointing with lbcrangebal
  4. DONE Rotator Center: Obtain images to measure the rotator center on chip 2. Chip2 is the only one to read out.
    1. Send a track preset at PA=0
    2. Send no preset (dec=-90) at PA=180,
    3. Send no preset at PA=90
    4. Send no preset at PA=270
  5. DONE Focal Plane Tilt Check: focal plane tilt with superfoc sequences at PA=0, PA=180, PA=0 /lbt/lbto/lbc/Calib_OBs/FOCUS/(OBNAME)
  6. DONE Observe with guiding and dithering: There are a few science-like OBs in the checkout OT
  7. DONE Observe a non-sidereal target: See https://wiki.lbto.org/Commissioning/NonSiderealGuiding#Details_for_sending_the_LBC_Prime_Focus_Non_45Sidereal_preset_in_TRACK_mode and https://scienceops.lbto.org/lbc/observing-with-lbc/non-sidereal-observing-with-the-lbcs/
    1. slow rates (rate*exptime <~ 2") with and without guiding
    2. faster rates (rate*exptime > 2") without guiding
  8. DONE Photometric Standards: If photometric, observe at least one photometric standard star field from each set (also tests dithering, but not guiding):
    1. Clem & Landolt
      • LBCB: UBV
      • LBCR: VRI
    2. sdssY
      • LBCB:gr (+Us as filler)
      • LBCR: rizY
  9. DONE TMS testing: Run TMS with 5,6,7. This will check:
    1. TMS operation and also
    2. also the new FPIA/TMS interaction.
  10. Pointing Model: Unavailable because of GPS
    1. plain LBC pointing model? or
    2. TMS pointing model?

MODS (high)

Overview
  1. DONE Rotator Zeropoints (Priority high)
  2. DONE Transform (Priority dependent on #3)
  3. DONE Spectrophotometric Standard (Priority high)
  4. DONE Photometric Standard field (Priority high)
  5. DONE Science-like observation (Priority low)
  6. Pointing model (Priority low)
Details
  1. DONE Rotator Zero Points
    • Scripts and MMS files are located: /home/lbto/Service_Observing/Restart_2023/MODS/RZP
    • These 3 masks cover the night. Try to align 2 masks. ~15 min each.
      • 20p25
      • 00p25
      • 04p25
  2. DONE Transform
    • MODS AGW Transform data - ~1" seeing - high (2hours to collect data on-sky, time during day to analyze, ~0.5 hr-1hr to verify on-sky) - must do RotZP First
    • Low priority if RotZP unchanged
    • Dither up and down a minimum of 15 cycles with pointing mask
    • Field and scripts located in MODS Restart OT and scripts located with Mask adjusted to pointing mask in:
      /nfs/192.168.39.20/homes/lbto/Service_Observing/Restart_2023/MODS/Transform
    • If RZP has changed:
      • Update transform from measurements
        • Files: agw.ini located on mods1 and mods2 computers in: /home/mods/Config/MODS1 or /home/mods/Config/MODS2
        • stop AGW services (mod1 stop agw, mods2 stop agw)
        • new file copied to ~/Config/MODSN/ and renamed agw.ini
        • Restart the AGW services (mods1 start agw, mods2 start agw)
      • Collect transform data for verification
      • Measure Rot ZP with MODS MOS masks
      • Dither Test M52 or M71
      • Confirmation new transform is acceptable
  3. DONE Spectrophotometric Standards
    • BD+33 until 04:30UT, BD+28 until 10UT, Feige110 6-10UT, G191-B2B from 9:30UT to twilight. Try to get Feige110 and/or G191-B2B which means standards are 2nd half of night.
    • If Clear, 2-3 spectra/mode
      For 3 spectra/mode, this will take 1hr/star
    • If not clear, 1 spectrum/mode
      One spectrum/mode will take 36 min
  4. DONE Photometric Standards
    • Available first half: Need to check nimgs in scripts (times based on scripts in Restart_2022).
      • M71 : twilight - 8UT: ~30 min (21 for the script and the extra for slew+collim)
        • dual (8 min)
        • red (7 min)
        • blue (6 min)
      • M92: twilight - 6 UT ~40 min
        • dual (12 min)
        • red (12 min)
        • blue (10 min)
      • M13: twilight - 5 UT
  5. DONE Dice 5 Science-like script
    • Similar to transform data collection so optional
  6. Pointing Model - 2hr low priority - Unavailable because of GPS

LUCIs (high)

Overview
  1. Rotator Zero Point (Priority high)
  2. Collect Transform Data (Priority high)
  3. Verify Transform Data (Priority high)
  4. Science-like seeing limited scripts (Priority med)
    1. Imaging - Sidereal tracking
    2. Imaging - Non sidereal tracking
    3. Longlist G210
    4. Longslit G200
  5. Science-like AO/ESM scripts (See AO checkout)
  6. Twilight Flats (Priority low)
  7. Photometric Standards (Priority low)
  8. Pointing Model (Priority Med)
Details
  1. Rotator Zero-Point (requires MOS mask) - (0.5 hr)
    • Masks:
      • ID 903033 - Stone O
      • ID 921934 - Stone A
      • iD 931027 - Stone D
    • To update results (if needed)
      • ssh -X into TO1 as telescope; then ssh - X as tcs@tcs1; The file location is in /home/telescope/TCS/Configuration/PCS/PCSInstrument.conf
    • Longslit Method (backup for no masks)
      • requires target field with two sources in slit, use alignment holes in longslit. Need LMS file with source in longslit (preferably two) - astrometric field ideal
  2. Collect transform data - (2hrs) - must do RotZP First
    • Target Fields, select one near transit:
      • Stone O 21 37 31 +00 31 55 Stone_O_m12.list
      • Stone P22 35 49 +01 31 33 Stone_P1.list
      • Stone A 00 54 02 -00 32 34 Stone_A3.list
    • Seeing should be no worse than ~1", list should include ~ 50 stars
    • Have the TO send an Active preset to tune pointing on the field and collimate
    • Configure LUCI with the N30 field stop aligned, N375 camera, H filter, flexure compensation on.
    • Follow the instructions as outlined on: https://wiki.lbto.org/Commissioning/ComTools#Guide_Probe_Transformation
    • Take 5xdarks to match exposure time taken
    • Process transform data (or shoot Doug an email ASAP)
  3. Verify Transform
    • Update transform
      • /lbt/oacontrol/current/etc/oacontrol.conf
      • rdwrconfig -u1 and rdwrconfig -u2
      • restart PCS and GCS
    • verify transform data (1 hr). The data collection process is the same as Transform data collection but a shortened list of 12-15 stars. Seeing should be no worse than 1.5"
  4. Science like scripts
    • Sidereal Imaging script
      • dither in field (0.5hr)
      • dither out of field (0.5hr)
    • non-sidereal Imaging Check
    • check longslit - high (1 hr)
      • Science-like OB for G210
      • Science-like OB for G200
  5. Twilight flats - low priority
    • N375 Imaging Flats (in order of priority)
      • K, H, J, z, Ks, BrGamma, FeII, H2, HeI, (Super low: J-low, J-high, P-beta, P-gamma, OH1060, OG1190, Y1, Y2)
    • N30 Imaging Flats
      • K, H, J, z, Ks
  6. Pointing Model - High (2hrs)
    • PtStarrs ~10th mag 12-20 stars
    • Poor seeing/clouds OK
  7. Photometric Standards - any seeing, must be clear
    • Scripts located in: Service_Observing/Restart_2023/LUCI/PhotStd

PEPSI (?? night -high)

Overview
  1. Throughput measurements
  2. Pointing Model
Details
  1. Throughput measurements (2 hrs over 3-4 nights)
  2. Pointing Model - 2hr Medium
    • Remove retro-reflecting dichroic (to give full FOV)
    • Take images with PFU (Active mode) -not a requirement, images come "free"
    • Need ~28 stars (10th mag or so)
      • PtStars (IRAF) will generate stars based on LST
      • minimized long azimuth slews
    • Analysis with TPOINT - D. Thompson
    • Can do in poor seeing/clouds

Mixed Mode Tests (0.5 night)

  • MODS/LUCI Mixed Mode
  • LBC/LUCI
  • LBC/MODS

LBTI AO and Science Cam (2x 0.5 night)

  • LBTI (2 x 0.5 night - need not be consecutive)
    • Night 1 (4.5hr):
      1. AO functional check-out on sky, close loop on a bright star, bayside stage positions for having the star on the science camera as well. 2h, both sides in parallel. Any Conditions.
      2. Science camera and operations check-out in parallel to above, plus 1h when the AO is done. Any Conditions.
      3. Find focus on both sides. 30 min per side. Requires good conditions
      4. Play with NCPA to see if we can improve it. 1h plus any extra time if we’re faster on above tasks. Requires good conditions
    • Night 2 (5hr):
      1. (pathlength correction issue prevents checkout) Find NOMIC nulling fringes and PhaseCam fringes, close the phase loop. 1h. Requires good conditions
      2. Find LMIRCam Fizeau fringes. 1h. Requires good conditions
      3. Try combination of M3 and roof mirror movement to rotate Fizeau fringes on LMIRCam, check/demonstrate that this has the intended effect. If time, try to align the fringes vertically on

LUCI AO (1 night)

  • System checkout
    • OSCO
    • Diffraction limited verification - use AO OT file - select a star, run through script
      • Bin 1 star (brighter than OSCO)
      • Bin 2 star
      • Bin 3 star
      • Bin 4 star
      • Bin 4 star faintest (WFSmag ~17.5-18)
      • Non-sideral (any brightness)
    • ESM verification (select a target from the ESM OT file, like M92, M52, etc)
      • Bin 1
      • Bin 2
      • Bin 4

Rainy Night Tasks

Support Files

MMS files:

Field MMS RA Dec MaskID Manufacture
Gaia 00+25 mods.2.g00d25ep.mms 00 00 00.139 +25 00 17.090 531713 2021
Gaia 20+25 mods.3.g2025ep2.mms 20 00 00.910 +25 00 28.100 531956 2021
Gaia 04+25 mods.4.g04d25ep.mms 03 59 58.669 +25 00 40.050 532732 2021

LMS Files:

Field RA (J2000) Dec (J2000) MASK ID
StoneO (new) 21 37 31.791 00 31 55.010 903033
StoneP (new) 22 35 49.616 01 31 33.950 933322
StoneA (new) 00 54 02.050 -00 32 34.460 921934

LMS Files with unboxed stars 1" longslit mask:

Field RA (J2000) Dec (J2000) Finder
M048stnOZP +975011 21 37 31.632 00 31 46.820 Stone O 1" longslit
M056stnPZP +912345 22 35 49.689 01 31 27.970 Stone P 1" longslit
M053stnAZP +991823 00 54 03.544 -00 32 12.120 Stone A 1" longslit
Scripts, LMS files, andfinder charts are located in the linked google drive. The OT RotZP _v2.xml contains the OB's used to generated the necessary scripts. Presets are sent with the normal LS_1.00arcsec slit mask. Rather than a longslit alignment, users should execute a MOS alignment, uploading manually the corresponding lms file. The RTD MOS mask alignment assumes that rather than slit #1 and #6 used across the top as is typically done, #1 is the central alignment hole in the bundle up top and #6 is the alignment hole in the bundle at the base of the mask. Once aligned, unboxed stars are used to refine the mask alignment. There are 2 stars located at the extremes of the slit that can be used for alignment verification. Andy questions, contact Jenny.

-- JenniferPower - 12 Jul 2023
Topic revision: r18 - 30 Oct 2023, JenniferPower
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback